# Solving rescheduling problems in heterogeneous urban railway networks using hybrid quantum-classical approach

### v3

@@ C. Disturbtions within the network and closures
In order to further evaluate the solutions, we present time-distance diagrams for network $7$, for the CPLEX solution and two different quantum realizations; these can be seen in Figs. 7,  8, and 9, respectively. We have chosen the line segment between GLC and KZ, offering a good demonstration. (We have introduced a small artificial ’distance’ between KO and KO(STM) wich are contiguous locations for sake of better visibility, hence the horizontal lines in the diagram.) The red dotted lines represent the conflicted situation, and the green solid lines represent the solution. All the three proposed solutions have similar figures of merit, they all provide a valid option to be realized. The conflicts are resolved in somewhat different ways, see e.g. the paths of trains 40150, 5312, and 26103. In details:
@@
This results in following objective ×$\times d_{\max}$ values: for CPLEX - $188.75$, for CQM first realization - $279$ for CQM second realization - $261.5$. Although, the first solution is optimal, it is reasonable to supply dispatchers with the range feasible solutions. Hence, these diagrams proved that CQM hybrid solver can produce usable solutions for railway traffic management.
@@
Figure 7: Real-life experiment. Time-distance diagram CPLEX solution of network $7$, see Tab. 2.
@@
Figure 8: Real-life experiment. Time-distance diagram CQM solution of network $7$ (see Tab. 2) first realization. Closures (yielding single track line) are between CB and KTC as well as between KO(STM) and KZ.
Figure 9: Real-life experiment. Time-distance diagram CQM solution of network $7$ (see Tab. 2) second realization.
7 Conclusions
We have demonstrated that quantum annealers can be readily applicable in train rescheduling optimization in urban railway networks. In particular we have encoded rescheduling problems in a linear integer program, and solved these using a state-of-the-art solver (CPLEX), as well as on the QA based CQM hybrid (quantum classical) solver. While the CQM hybrid solver does not outperform classical solvers in general, we have found specific cases in which they were actually better. This supports the expectation that future quantum devices will be efficient in solving large scale-problems, and in our case, real-time railway dispatching problems on the national scale; even in the range that are beyond the scope of current exact models and heuristics.
While the quantum-based solvers would possibly return suboptimal solutions, these are still often better than the solutions obtained manually or based on smaller scale models. More importantly, all generated solutions were feasible. In each computation, the CQM hybrid solver always has taken some advantage of the QPU. Hence, we believe that such a little help from the QPU must have boosted the classical heuristics in the hybrid solver.
Recall that quantum annealer devices are subjects of many ongoing debates. Previous studies claimed that the quantum nature of their operation is limited [45], while recent works argue that their operation relies more on the thermalization with a cold environment in place of actual quantum adiabatic evolution [6]. They are often criticized for the small size of problems they can address and various classical solvers do outperform them on certain problems at the present state of the art. In spite of all these, our computational results demonstrate the readiness of hybrid (quantum classical) solvers to handle real-life railway problems. We believe that future investigations shall focus on hybrid approaches both for railway applications as well as more generally.
Finally, several future research directions can be determined. First concerns the ongoing worldwide works on the development of fault tolerant quantum computing devices [17] with proper error correction/mitigation, this will boost results of approaches supported by quantum computing (and QA in particular). Second concerns creation of custom open source hybrid solver that is dedicated for railway problems and will use noise of quantum device to our advantage e.g. by modelling some stochastic behaviour on railways. Third concerns evolution and application of quantum inspired techniques such as tensor networks technique, simulated bifurcation [26]. With all these future developments, QA-based approaches tend to become more powerful and ready to address for the first time large-scale real-time challenges in railway traffic management, which are currently unsolvable with traditional techniques.
Data availability
Acknowledgements
This research was supported by the Ministry of Culture and Innovation and the National Research, Development and Innovation Office within the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004) (MK), and by the Silesian University of Technology Rector’s Grant no. BKM - 720/RT2/2023 12/020/BKM_2023/0252 (KK). This research was funded in part by National Science Center, Poland, under grant number 2019/33/B/ST6/0201 (LB) and 2023/07/X/ST6/00396 (KD). For the purpose of Open Access, the authors has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
We would like to thank to Sebastian Deffner as well as to the whole scientific community of the Department of Physics, University of Maryland, Baltimore County (UMBC) for valuable discussion; and to Bartłomiej Gardas, and Zbigniew Puchała for valuable motivation; and to Katarzyna Gawlak, Akash Kundu, and Özlem Salehi for data validation and assistance with coding. We acknowledge the cooperation with Koleje Ślaskie sp. z o.o. (eng. Silesian Railways).
Compliance with ethical standards
Conflict of interest
The authors declare that they have no conflict of interest.
Ethical approval
This article does not contain any studies with human participants or animals performed by any of the authors.
Appendix A. Comparative analyses for different values of the t_min parameter on selected solutions
The right choice of the t_min parameter can improve results meaningfully. Hence, in this Appendix we analyse in more detail the role of the t_min solver parameter for selected problems in Section 6.3. In particular, for networks $7$ and $9$ therein, we have performed runs with t_min sweeping over a range $[5,60]$; the results are presented in Fig. 10 and Fig. 11. The objective value is roughly log-linear in computational time. For small t_min parameter values the computational time is shorter, whereas for large t_min it grows exponentially. Overall, the improvement in the objective is overwhelmed by the increase of the computational time. This is important from the point of view of the algorithm layout, as it may be tied to a power law scaling. Such a behaviour is plausible in the case of QA methods [46, 1, 20].
Figure 10: Real-life experiments. The sweep over t_min parameter for network $7$, see Tab. 2. All solutions were feasible. For higher t_min we expect smaller objective values but higher total computational time. For each parameter value there were $5$ realizations of the experiment were performed. Interestingly, a linear (negatively sloped) relation between the objective and the logarithm of ${\tt t\_min}$ can be observed, suggesting power law scaling.
Figure 11: Real-life experiments. The sweep over t_min parameter for network $9$, see Tab. 2. All solutions were feasible, results were similar to those in Fig. 10.

### v4

@@ Appendix A. Train diagrams of CPLEX and CQM
To further evaluate the solutions, we present time-distance diagrams for network $7$, for the CPLEX solution and two different quantum realizations; these can be seen in Fig. 6 (left CPLEX, middle and right two realizations of CQM). We have chosen the line segment between GLC and KZ, offering a good demonstration. (We have introduced a small artificial ’distance’ between KO and KO(STM) which are contiguous locations for sake of better visibility, hence the horizontal lines in the diagram.) The red dotted lines represent the conflicted situation, and the green solid lines represent the solution. All the three proposed solutions have similar figures of merit, they all provide a valid option to be realized. The conflicts are resolved in somewhat different ways, see e.g. the paths of trains 40150, 5312, and 26103. In details:
@@
This results in the following objective ×$\times d_{\max}$ values: for CPLEX - $188.75$, for CQM first realization - $279$ for CQM second realization - $261.5$. Although the first solution is optimal, it is reasonable to supply dispatchers with the range of feasible solutions. Hence, these diagrams proved that the CQM hybrid solver can produce usable solutions for railway traffic management.
@@
Figure 6: Real-life experiment. Time-distance diagram solution of network $7$ CPLEX (left) and two realizations of CQM (middle and right), see Tab. 2. Real-life experiment. Time-distance diagram CQM solution of network $7$ (see Tab. 2) first realization. Closures (yielding single-track line) are between CB and KTC as well as between KO(STM) and KZ.